A sample of 10 -2 kg each of 2 substances A and B with half lives of 4 and 8 sec respectively. The atomic weight of A and B are in the ratio of 1: 2. Find the amount of A&B after an interval of 16 sec. Also the report the ratio of initial rates of decay these two.

A sample of 10-2 kg each of 2 substances A and B  with half lives of 4 and 8 sec respectively.  The atomic weight of A and B are in the ratio of 1: 2.  Find the amount of A&B after an interval of 16 sec.  Also the report the ratio of initial rates of decay these two.

Grade:11

1 Answers

askiitian.expert- chandra sekhar
14 years ago

Hi Kiran,

atomic weights ratio of A and B=1:2

implies the no of molecules of B in the sample will be half the no of molecules of A

after 16 seconds:

A will be 10-2/16 kg

B will be 10-2/4 kg

ratio of initial rates = k1*[A]/k2*[B]

k1/k2=8/4 =2  and [A]/[B] = 2

ratio of initial rates = k1*[A]/k2*[B] = 2*2=4